Set up a system of three equations with three variables.
A, B, and C complete a task in 3 days. On the first day they do 1/3 of the task with A working 10 hours, B working 4 hours, and C working 10 hours. On the second day they do 1/4 of the task with A working 4 hours, B working 6 hours, and C working 9 hours. They finish the task on the third day with A working 7 hours, B working 12 hours, and C working 12 hours. Find the number of hours it would take for each man working alone to do the task.

Answer by richwmiller(9135)

10a+4b+10c=1/3;
4a+6b+9c=1/4;
7a+12b+12c=1-1/3-1/4
a = 1/60, b = 1/72, c = 1/90
it would take a 60 days working alone
it would take b 72 days working alone
it would take c 90 days working alone